OpenStudy (anonymous):

f(x)= (1.6^3)*sqrt(x) and g(x)= x^2. find intersection points.

OpenStudy (campbell_st):

equate the functions \[1.6^3\sqrt{x} = x^2\] divide both sides by root x \[1.6^3 = \frac{x^2}{x^{1/2}}\] using index laws \[1/6^3 = x^{3/2}\] \[x = (1.6^3)^{2/3}\] \[x = 1.6^2\]

OpenStudy (anonymous):

from the second step: for division with like bases, would subtract 2 - 1/2?

OpenStudy (campbell_st):

you need to substitute the x value to find y y = x^2 then y = (1.6^2)^2 y = 1.6^4 the ordered pair is (1.6^2, 1.6^4)

OpenStudy (campbell_st):

and there is another point of intersection , I just realised... when x=0, then y= 0 so the points of intersection and (0,0) and (1.6^2, 1.6^4)

OpenStudy (anonymous):

what about rotating the region about the y-axis and finding resulting volume using shell method?

OpenStudy (campbell_st):

I don't know the shell method

OpenStudy (anonymous):

\[2pi \int\limits_{a}^{b}(f(x)-g(x))dx\]

OpenStudy (campbell_st):

all I use for volume of a solid revolved is \[V = \pi \int\limits_{a}^{b} y^2 dx \]

OpenStudy (campbell_st):

well the problem is \[V = \pi \int\limits_{0}^{1.6^2} 1.6^3x^{1/2} - x^2 dx\] rough guess its \[V = \pi[ 1.6^3 2/3x^3/2 - 1/3x^3]_{0}^{1.6^2}\]

OpenStudy (campbell_st):

oops \[V = \pi[1.6^3\frac{2}{3}x^{2/3} - \frac{1}{3}x^3]_{0}^{1.6^2}\]

OpenStudy (campbell_st):

just never heard it referred to as the shell method

OpenStudy (anonymous):

the one you are using is called disc/washer method

OpenStudy (campbell_st):

ok... thats all I use sorry I think the integral is correct... you just need a multiple of 2 out the front with pi

OpenStudy (campbell_st):

f(x) is the upper function and g(x) is the lower function.

OpenStudy (anonymous):

thank you for your help. :-)

OpenStudy (anonymous):

\[2pi \int\limits_{a}^{b}x(f(x)-g(x))dx\] that is the shell method, sorry. forgot the x.

OpenStudy (anonymous):

have a nice day/night.